OpenStudy (adfksfnads):

Name the different forms of linear equations you learned. If (1,2) and (3,4) lie on the same line, give the slope-intercept equation for the line.

OpenStudy (dumbcow):

\[y = m(x-x_1) + y_1\] where \[m = \frac{y_2 - y_1}{x_2 - x_1}\]

TheSmartOne (thesmartone):

There are three forms, you can read through this to learn more about it :) http://www.mathsisfun.com/algebra/linear-equations.html

OpenStudy (adfksfnads):

I have answered the first part of the question. I just need help with the second part.

OpenStudy (adfksfnads):

@agent0smith , just the second part

OpenStudy (agent0smith):

I don't know what the second part is. Just post a brand new question clearly showing what you need help with.